• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 15:57:47 on March 5,2001

    Entry number 57777

    keyword=Shift Summary

    Shift Summary

    0800 Old Coin Kin_6 starting.
    0835-0844 Run 2525, LD2 (rad. in) 1M events
    0844-0854 Run 2526, LD2 (rad. in) 1M events
    0855-0903 Run 2527, LD2 (rad. in) 1M events
    0905-0913 Run 2528, LD2 (rad. out) 350 k events
    0914-0930 Run 2529, LD2 (rad. out) 1M events
    Hall A Going to Restricted access.
    Changing to He target
    Fixing eP target tape.
    Fix Right arm angle camera.
    Survey at present angles. Left arm = 37.46 deg, Right arm = 49.98 deg
    Move spectrometer to Coin_6 settings:
    Left arm = 37.37 deg, Right arm = 49.73 deg
    0945: P_Left = 1.2766 GeV/c, P_Right = -1.015 GeV/c
    1600: Hall is still in restricted access